# Tree-Planted Transformers: Large Language Models with Implicit Syntactic Supervision

### v1

@@ 4.1 Settings
We used LG dataset of Hu et al. (2020), which comprises approximately 42M tokens from BLLIP corpus (Charniak et al., 2000). Implicit syntactic supervision with each of three types of syntactic structures was investigated: (i) dependency structures ([dep.]), (ii) constituency structures ([cons.]), and (iii) binarized constituency structures ([bin.]). The (i) dependency structures were parsed with the en_core_web_sm model from the spacy library (Montani et al., 2023).777https://spacy.io The (ii) constituency structures were re-parsed with the Berkeley Neural Parser (Kitaev and Klein, 2018)888https://github.com/nikitakit/self-attentive-parser by Hu et al. (2020). The (iii) binarized constituency structures were obtained by the binarization of the (ii) constituency structures with the chomsky_normal_form function from the nltk library (Bird et al., 2009).999https://www.nltk.org
@@
We used the same architecture and BPE tokenizer as GPT-2 small (124M; Radford et al., 2018). The implementation of GPT2LMHeadModel and GPT2Tokenizer from the transformers library (Wolf et al., 2020)101010https://huggingface.co/docs/transformers were employed but all parameters of GPT2LMHeadModel were randomly initialized. For the tree-planted head and the weight of the tree-planting loss, we adopted a single attention head on the last layer and , respectively. The choice of the tree-planted head and the weight was based on preliminary experiments and the detailed effects of them will be described in Section 5.
@@ 4.2 Overall accuracies
Table 3 shows the overall accuracies of TPTs and their baselines on the SyntaxGym benchmark (SG), along with word-level perplexity on the BLLIP test set (PPL). The overall accuracies were calculated across the syntactic suites. We also report the accuracies of several SLMs that were also trained on the same BLLIP-LG dataset: PLM, PLM-mask (Qian et al., 2021), and TG (Sartran et al., 2022). Only unmarked PLM and PLM-mask can be fairly comparable with TPTs as their evaluation was conducted generating a single syntactic structure via greedy search, to align inference costs with TPTs.111111The fair comparison of TG was not performed because their trained parameters were not publicly available. and represent the reference points from Sartran et al. (2022) as their inference methods are more costly than TPTs: and employed word-synchronous beam search (Stern et al., 2017) of action beam size 100121212Word beam size was 10 and fast track size was 5. and the external parser (Dyer et al., 2016), respectively. The accuracies of several LLM-like models are also reported from Sartran et al. (2022): GPT-2 (Radford et al., 2018), Gopher (Rae et al., 2022), and Chinchilla (Hoffmann et al., 2022). They are also the reference points as these LLM-like models were trained on significantly larger corpora (denoted by ). Perplexity can be directly comparable only within the same mark, either or , due to differences in the tokenization of the constituency parser and dependency parser.
@@
TPT[zero], which is equivalent to a Transformer LM, underperformed all TPTs with some implicit syntactic supervision, suggesting that tree-planting can introduce syntactic knowledge to TPTs.
@@
Among TPTs with some implicit syntactic supervision, TPT[dep.] achieved the best performance. We further investigate this point in Subsection 4.3.
Most importantly, despite the lack of explicit syntactic supervision, TPTs[dep.][cons.] significantly outperformed not only the comparable SLMs (unmarked PLM and PLM-mask) but also the various SLMs that generate hundreds of syntactic structures in parallel (PLM and PLM-mask).
Even though the best TPT[dep.] underperformed the reference points of the more costly TG and the larger LLM-like models, these observations adequately suggest that tree-planting and TPTs are the promising foundation for SLLMs.
@@ 4.3 Circuit accuracies
In this subsection, we investigate the advantages of dependency structures through the lens of circuit accuracies. Figure 3 shows the circuit accuracies of TPTs with some implicit syntactic supervision and the baseline model with zero weight for the tree-planting loss on the SyntaxGym benchmark. The circuit accuracies calculated across the syntactic suites (the vertical axis) are plotted against the models (the horizontal axis), with each dot denoting the accuracy of a specific seed.
@@
TPT[dep.] outperformed TPT[zero] on 5/6 circuits, suggesting that syntactic supervision of dependency structures is generally advantageous over zero supervision. However, the Garden-Path Effects circuit presents an exception, where LMs are evaluated for the ability to be surprised in a human-like manner, through comparisons between sentences minimally different not in grammaticality but in local ambiguity (Hu et al., 2020). The underperformance of TPT[dep.] may suggest that due to the syntactic knowledge introduced by tree-planting with dependency structures, TPT[dep.] was no longer surprised by locally ambiguous but grammatical sentences. We further investigate this point in Appendix B.
@@
TPT[dep.] outperformed TPT[bin.] on 3/6 circuits, with similar performance (a difference less than ) on the other 3 circuits. Notably, TPT[dep.] achieved significantly better performance (a difference more than ) on the Agreement and Licensing circuits. Noji and Oseki (2021) reported that deep syntactic supervision is not always optimal; rather mild syntactic supervision is sufficient for addressing long-distance dependencies between elements within and outside complex NP subjects. Given that (i) the Agreement and Licensing circuits consist only of syntactic suites that exemplify this condition131313Among the other syntactic circuits, the Center Embedding circuit also exemplifies this condition. and (ii) the average syntactic distance in the training data is significantly shorter for dependency structures (4.8) than binarized constituency structures (13.1), it could be argued that dependency structures would be more suitable as “good enough” syntactic supervision than binarized constituency structures.141414The average syntactic distance of constituency structures is 10.0. This suggests that dependency structure would also be superior to constituency structure as “good enough” syntactic supervision, besides the points discussed in the “vs. constituency structures” paragraph.
@@ 5.1 Number of tree-planted heads
Our TPTs are based on a 12-layer, 12-head Transformer LM. In Section 4, out of heads, we adopted a single attention head on the last layer as a tree-planted head. In this subsection, we explore two alternatives: (i) head-direction extension and (ii) layer-direction extension. For the head-direction extension, 0, 1, 3, 6, 9, and 12 heads on the last layer were adopted as tree-planted heads. For the layer-direction extension, one attention head from the each of bottom 0, 1, 3, 6, 9, and 12 layers was adopted as tree-planted heads.
@@
Considering the overall accuracies on the SyntaxGym benchmark, in both the head-direction and layer-direction extension, the highest accuracy was achieved when only a single head was adopted as a tree-planted head, while it is noteworthy that all the models with tree-planted heads outperformed the model without them. Incidentally, it should be mentioned that the result of the layer-direction extension exhibited significantly more variability. Although the reason why a single tree-planted head would work well is unclear, the adoption of multi tree-planted heads inherently induces the handling of redundant information across heads, which might potentially hinder the management of non-syntactic information of natural languages (e.g., lexical information). Regarding perplexity, no consistent trend emerged.
@@ 5.2 Weight of a tree-planting loss
In Section 4, we adopted as the weight of the tree-planting loss. Here, we extend to , , , , and . are equivalent to TPT[zero] and TPT[dep.], respectively.
The rightmost column of Figure 4 shows the results of the weight extension. The overall accuracies on the SyntaxGym benchmark display a single-peaked pattern, with the maximum reached for . Interestingly, this result suggests that by overtly focusing on reflecting syntactic structures, TPTs paradoxically become unable to learn syntactic knowledge. On the other hand, we observed that the perplexity got worse monotonically as the weight increased. From these observations, we may deduce that to acquire syntactic knowledge, TPTs should learn not only to reflect syntactic structures in their attention weights but also to precisely predict the next word. Therefore, the weight of the tree-planting loss emerges as a critical hyperparameter, indicating that the search for the optimal balance between the next-word prediction loss and tree-planting loss is vital for developing more human-like TPTs.
@@ 6 Conclusion
In this paper, we proposed a novel method dubbed tree-planting: implicitly “plant” trees into attention weights of Transformer LMs to reflect syntactic structures of natural language. Specifically, Transformer LMs trained with tree-planting are called Tree-Planted Transformers (TPT), which learn syntax on small treebanks via tree-planting and then scale on large text corpora via continual learning with syntactic scaffolding. Targeted syntactic evaluations on the SyntaxGym benchmark demonstrated that TPTs, despite the lack of explicit syntactic supervision, significantly outperformed various SLMs with explicit syntactic supervision that generate hundreds of syntactic structures in parallel, suggesting that tree-planting and TPTs are the promising foundation for SLLMs.
@@
There are at least three limitations in this paper. First, we only conducted sentence-level tree-planting. Typically, LLMs are trained at the document level, but SLMs are trained at the sentence level (Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021; Yoshida and Oseki, 2022), because on treebanks, the annotations are assigned at the sentence level. Because of this constraint, we also employed sentence-level experimental design and verified the effectiveness of the proposed method first and foremost. Recent research in SLMs, however, has begun to extend treebank annotations to the document level and train document-level SLMs on them (Sartran et al., 2022; Murty et al., 2023). When constructing TPTs for practical use, it might be beneficial to follow these recent studies and perform tree-planting with document-level annotations.
Second, we only evaluated TPTs on the syntactic knowledge benchmark and perplexity. Recently, Murty et al. (2023) evaluated the performance of SLMs on tasks other than the targeted syntactic evaluations for the first time, suggesting that syntactic knowledge could also be beneficial to solving them. This indicates that there is also room for a broader evaluation of our methodology.
Finally, the development of a novel continual learning method (e.g., updating the parameters of tree-planted heads sparingly) would be necessary for scaling TPTs on large corpora, without compromising the syntactic knowledge but rather exploiting it as syntactic scaffolding. In future work, we plan to develop a novel method for "climbing trees" in TPTs.
@@
A significant feature of TPT lies in the training efficiency, which can potentially contribute to reducing computational resources. One minor concern is the possibility of bias in the models utilized in this paper, attributed to the training data (i.e., the BLLIP corpus), although this experimental setting follows conventional practices in the literature on SLMs. We employed ChatGPT and Grammarly for writing assistance, and for the development of experimental code, we utilized ChatGPT and Copilot. These tools were used in compliance with the ACL 2023 Policy on the Use of AI Writing Assistance.
@@ Acknowledgements
We would like to thank Peng Qian for providing us with the re-parsed BLLIP-LG dataset, which is used to train TPT[cons.]. We also wish to thank Laurent Sartran for answering various questions. We are also grateful to Kohei Kajikawa, Shinnosuke Isono, Yushi Sugimito, and Taketeru Yamakoshi for their insightful comments and suggestions. This work was supported by JST PRESTO Grant Number JPMJPR21C2.
@@ Appendix B Further investigation of the Garden-Path Effects circuit
In Subsection 4.3, we suggest the probability that syntactic knowledge introduced by tree-planting with dependency structures may prevent TPT[dep.] from being surprised by locally ambiguous but grammatical sentences. To inspect this, we break down the Garden-Path Effects circuit into the syntactic suites: “main verb / reduced relative clause” (MVRR) and “NP/Z garden-paths” (NP/Z).
@@
Conversely, TPT[cons.][bin.] did not underperform TPT(zero.) on the MVRR suites. This result could be straightforwardly understood, given that on these structures, the participle verb (seen/walked) and the head of the subject NP (dog) are equidistant from the main verb (chased). However, it is worth noting that the determiner of the subject NP (the) also shares this distance, which may not always be a desirable property (cf. Subsection 4.3).
Finally, PLM-mask, the more costly SLM, also underperformed TPT[zero] on the MVRR suites. This suggests that the models with explicit syntactic supervision may also struggle with losing sensitivity to the local ambiguity as PLM[dep.].
@@ Appendix D Begin/End Of Sentence Tokens
For constituency structures, we introduced the BOS/EOS tokens by modifying the tree structure to encapsulate the original structure within a new root node, specifically by adding a BOS token and an EOS token as the first and the last child of this new root, respectively.
@@ Appendix E License of the data/tools
Data/tool |
License |
BLLIP (Charniak et al., 2000) |
BLLIP 1987-89 WSJ Corpus Release 1 License Agreement |
SyntaxGym (Gauthier et al., 2020) |
MIT |
spacy (Montani et al., 2023) |
MIT |
nltk (Bird et al., 2009) |
Apache 2.0 |
transformers (Wolf et al., 2020) |
Apache 2.0 |
Berkeley Neural Parser (Kitaev and Klein, 2018) |
MIT |
PLM/PLM-mask (Qian et al., 2021) |
MIT |

### v2

@@ 4.1 Settings
We used LG dataset of Hu et al. (2020), which comprises approximately 1.8M sentences from BLLIP corpus (Charniak et al., 2000). Implicit syntactic supervision with each of three types of syntactic structures was investigated: (i) dependency structures ([dep.]), (ii) constituency structures ([cons.]), and (iii) binarized constituency structures ([bin.]). The (i) dependency structures were parsed with the en_core_web_sm model from the spacy library (Montani et al., 2023).999https://spacy.io The (ii) constituency structures were re-parsed with the Berkeley Neural Parser (Kitaev and Klein, 2018)101010https://github.com/nikitakit/self-attentive-parser by Hu et al. (2020). The (iii) binarized constituency structures were obtained by the binarization of the (ii) constituency structures with the chomsky_normal_form function from the nltk library (Bird et al., 2009).111111https://www.nltk.org We removed 43,986 sentences that the dependency parser analyzed as multiple sentences, but the constituency parser analyzed as a single sentence.
@@
We used the same architecture and BPE tokenizer as GPT-2 small (124M; Radford et al., 2018). The implementation of GPT2LMHeadModel and GPT2Tokenizer from the transformers library (Wolf et al., 2020)121212https://huggingface.co/docs/transformers were employed but all parameters of GPT2LMHeadModel were randomly initialized. For the tree-planted head and the weight of the tree-planting loss, we adopted a single attention head on the last layer and , respectively. The choice of the tree-planted head and the weight was based on preliminary experiments and the detailed effects of them will be described in §5.
@@ 4.2 Overall accuracies
Table 2 shows the overall accuracies of TPTs and their baselines on the SyntaxGym benchmark (SG), along with word-level perplexity on the BLLIP test set (PPL). The overall accuracies were calculated across the syntactic suites. We also report the accuracies of several SLMs that were also trained on the same BLLIP-LG dataset: PLM, PLM-mask (Qian et al., 2021), and TG (Sartran et al., 2022). Only unmarked PLM and PLM-mask can be fairly comparable with TPTs as their evaluation was conducted generating a single syntactic structure via greedy search, to align inference costs with TPTs.131313The fair comparison of TG was not performed because their trained parameters were not publicly available. and represent the reference points from Sartran et al. (2022) as their inference methods are more costly than TPTs: and employed word-synchronous beam search (Stern et al., 2017) of action beam size 100141414Word beam size was 10 and fast track size was 5. and the external parser (Dyer et al., 2016) to generate 300 candidate structures, respectively. The accuracies of several LLM-like models are also reported from Sartran et al. (2022): GPT-2 (Radford et al., 2018), Gopher (Rae et al., 2022), and Chinchilla (Hoffmann et al., 2022). They are also the reference points as these LLM-like models were trained on to larger corpora (denoted by ). Perplexity can be directly comparable only within the same mark, either or , due to differences in the tokenization of the dependency parser and constituency parser.
@@
TPT[zero], which is equivalent to a Transformer LM, underperformed all TPTs with some implicit syntactic supervision, suggesting that tree-planting can induce data-efficient syntactic generalization.
@@
Among TPTs with some implicit syntactic supervision, TPT[dep.] achieved the best performance. We further investigate this point in §4.3.
Most importantly, despite the lack of explicit generation of syntactic structures, TPTs[dep.][cons.] significantly outperformed not only the comparable SLMs (unmarked PLM and PLM-mask) but also the various SLMs that generate hundreds of syntactic structures in parallel (PLM and PLM-mask).
Even though the best TPT[dep.] underperformed the reference points of TG, which consumed at least higher inference cost, and LLM-like models, which were trained on at least larger data, these observations adequately suggest that TPTs can learn human-like syntactic knowledge as data-efficiently as SLMs while maintaining the modeling space of Transformer LMs unchanged.
@@ 4.3 Circuit accuracies
In this subsection, we investigate the reason for the high compatibility between tree-planting and dependency structures through the lens of circuit accuracies. Figure 3 shows the circuit accuracies of TPTs with some implicit syntactic supervision and the baseline model with zero weight for the tree-planting loss on the SyntaxGym benchmark. The circuit accuracies calculated across the syntactic suites (the vertical axis) are plotted against the models (the horizontal axis), with each dot denoting the accuracy of a specific seed.
@@
TPT[dep.] outperformed TPT[zero] on 5/6 circuits, suggesting that tree-planting with dependency structures is generally advantageous over zero supervision. However, the Garden-Path Effects circuit presents an exception, where LMs are evaluated for the ability to be surprised in a human-like manner, through comparisons between sentences minimally different not in grammaticality but in local ambiguity (Hu et al., 2020). The underperformance of TPT[dep.] may suggest that due to the syntactic knowledge introduced by tree-planting with dependency structures, TPT[dep.] was no longer surprised by locally ambiguous but grammatical sentences. We further investigate this point in Appendix B.
@@
TPT[dep.] outperformed TPT[bin.] on 3/6 circuits, with similar performance (a difference less than ) on the other 3 circuits. Notably, TPT[dep.] achieved significantly better performance (a difference more than ) on the Agreement and Licensing circuits. Noji and Oseki (2023) reported that deep syntactic supervision is not always optimal; rather mild syntactic supervision is sufficient for addressing long-distance dependencies between elements within and outside complex NP subjects. Given that (i) the Agreement and Licensing circuits consist only of syntactic suites that exemplify this condition151515Among the other syntactic circuits, the Center Embedding circuit also consist only of syntactic suites that exemplify this condition. and (ii) the average syntactic distance in the training data is significantly shorter for dependency structures (4.8) than binarized constituency structures (13.1), it could be argued that tree-planting would be more “good enough” syntactic supervision with dependency structures, rather than with binarized constituency structures.161616The average syntactic distance of constituency structures is 10.0. This suggests that tree-planting would also be more “good enough” syntactic supervision with dependency structures rather than with constituency structures, besides the points discussed in the “vs. constituency structures” paragraph.
@@ 5.1 Number of tree-planted heads
Our TPTs are based on a 12-layer, 12-head Transformer LM. In §4, out of heads, we adopted a single attention head on the last layer as a tree-planted head. In this subsection, we explore two alternatives: (i) head-direction extension and (ii) layer-direction extension. For the head-direction extension, heads on the last layer were adopted as tree-planted heads. For the layer-direction extension, one attention head from each of the bottom layers was adopted as tree-planted heads.
@@
Considering the overall accuracies on the SyntaxGym benchmark, in both the head-direction and layer-direction extension, the highest accuracy was achieved when only a single head was adopted as a tree-planted head, while it is noteworthy that all the models with tree-planted heads outperformed the model without them. Incidentally, it should be mentioned that the result of the layer-direction extension exhibited significantly more variability. Although the exact reason why a single tree-planted head would work well is unclear, the adoption of multi tree-planted heads inherently induces the handling of redundant information across heads, which might potentially hinder the management of non-syntactic information of natural languages (e.g., lexical information). Regarding perplexity, no consistent trend emerged.
@@ 5.2 Weight of a tree-planting loss
In §4, we adopted as the weight of the tree-planting loss. Here, we extend to , , , , and . are equivalent to TPT[zero] and TPT[dep.], respectively.
The rightmost column of Figure 4 shows the results of the weight extension. The overall accuracies on the SyntaxGym benchmark display a single-peaked pattern, with the maximum reached for . Interestingly, this result suggests that by overtly focusing on reflecting syntactic structures, TPTs paradoxically become unable to learn syntactic knowledge efficiently. On the other hand, we observed that the perplexity got worse monotonically as the weight increased. From these observations, we may deduce that to acquire syntactic knowledge efficiently, TPTs should learn not only to reflect syntactic structures in their attention weights but also to precisely predict the next word. Therefore, the weight of the tree-planting loss emerges as a critical hyperparameter, indicating that the search for the optimal balance between the next-word prediction loss and tree-planting loss is vital for developing more human-like TPTs.
@@ 6 Conclusion
In this paper, we propose a new method dubbed tree-planting: instead of explicitly generating syntactic structures, we “plant” trees into attention weights of unidirectional Transformer LMs to implicitly reflect syntactic structures of natural language. Specifically, unidirectional Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which inherit the training efficiency from SLMs without changing the inference efficiency of their underlying Transformer LMs. Targeted syntactic evaluations on the SyntaxGym benchmark demonstrated that TPTs, despite the lack of explicit generation of syntactic structures, significantly outperformed not only vanilla Transformer LMs but also various SLMs that generate hundreds of syntactic structures in parallel. This result suggests that TPTs can learn human-like syntactic knowledge as data-efficiently as SLMs while maintaining the modeling space of Transformer LMs unchanged.
@@
This paper has at least three limitations. First, we only conducted sentence-level tree-planting. Typically, Transformer LMs are trained at the document level, but SLMs are trained at the sentence level (Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021; Yoshida and Oseki, 2022), because on treebanks the annotations are assigned at the sentence level. Because of this constraint, we also employed sentence-level experimental design and verified the effectiveness of the proposed method first and foremost. Recent research in SLMs, however, has begun to extend treebank annotations to the document level and train document-level SLMs on them (Sartran et al., 2022; Murty et al., 2023). When constructing TPTs for practical use, it might be beneficial to follow these recent studies and perform tree-planting with document-level annotations.
Second, we only evaluated TPTs on the syntactic knowledge benchmark and perplexity. Recently, Murty et al. (2023) evaluated the performance of SLMs on text classification tasks; to the best of our knowledge, this is the first work that evaluated SLMs on tasks other than the targeted syntactic evaluations. More recently, Hu et al. (2024) evaluated unsupervised SLMs on text generation tasks as well as text classification tasks. Murty et al. (2023) and Hu et al. (2024) both suggested that syntactic supervision could also be beneficial to solving them; this indicates that there is also room for a broader evaluation of our methodology.
Finally, there might still be room for further improvement of tree-planting and TPTs. For instance, an in-depth comparison between TPT and TG, which achieved much better accuracy on the Center Embedding and Garden-Path Effects circuits (cf. Sartran et al., 2022), could provide insights for improving the design of tree-planting. Additionally, since the modeling space of TPT is identical to that of underlying Transformer LMs, TPTs are theoretically capable of continual learning on standard text corpora. In future work, we plan to develop a novel method to scale TPTs on large text corpora without compromising syntactic knowledge.
@@
A significant feature of TPT lies in the training efficiency and inference efficiency, which can potentially contribute to reducing computational resources. One minor concern is the possibility of bias in the models utilized in this paper, attributed to the training data (i.e., the BLLIP corpus), although this experimental setting follows conventional practices in the literature on SLMs. We employed ChatGPT and Grammarly for writing assistance and utilized ChatGPT and Copilot for the development of experimental code. These tools were used in compliance with the ACL 2023 Policy on the Use of AI Writing Assistance.
@@ Acknowledgements
We appreciate the insightful reviews provided by the three anonymous ARR reviewers. We would also like to thank Peng Qian for supplying the re-parsed BLLIP-LG dataset, which was used to train TPT[cons.]. Special thanks to Laurent Sartran for answering various questions regarding Sartran et al. (2022). We are grateful to Kohei Kajikawa, Shinnosuke Isono, Yushi Sugimito, and Taketeru Yamakoshi for their valuable comments and suggestions. This work was supported by JSPS KAKENHI Grant Number 24H00087, Grant-in-Aid for JSPS Fellows JP24KJ0800, JST PRESTO Grant Number JPMJPR21C2, and JST SPRING Grant Number JPMJSP2108.
@@ Appendix B Further investigation of the Garden-Path Effects circuit
In §4.3, we suggest the probability that syntactic knowledge introduced by tree-planting with dependency structures may prevent TPT[dep.] from being surprised by locally ambiguous but grammatical sentences. To inspect this, we break down the Garden-Path Effects circuit into the syntactic suites: “main verb / reduced relative clause” (MVRR) and “NP/Z garden-paths” (NP/Z).
@@
Conversely, TPT[cons.][bin.] did not underperform TPT[zero.] on the MVRR suites. This result could be straightforwardly understood, given that on these structures, the participle verb (seen/walked) and the head of the subject NP (dog) are equidistant from the main verb (chased). However, it is worth noting that the determiner of the subject NP (the) also shares this distance, which may not always be a desirable property for tree-planting (cf. §4.3).
Finally, PLM-mask, the more inference-costly SLM, also underperformed TPT[zero] on the MVRR suites. This suggests that the models with explicit syntactic supervision may also struggle with losing sensitivity to the local ambiguity as PLM[dep.].
@@ Appendix D Begin/End Of Sentence Tokens
For constituency structures, we introduced BOS/EOS tokens by modifying the tree structure to encapsulate the original structure within a new root node, specifically by adding a BOS token and an EOS token as the first and the last child of this new root, respectively.
@@ Appendix E License of the data/tools
| Data/tool | License |
| BLLIP (Charniak et al., 2000) | BLLIP 1987-89 WSJ Corpus Release 1 License Agreement |
| SyntaxGym (Gauthier et al., 2020) | MIT |
| spacy (Montani et al., 2023) | MIT |
| nltk (Bird et al., 2009) | Apache 2.0 |
| transformers (Wolf et al., 2020) | Apache 2.0 |
| Berkeley Neural Parser (Kitaev and Klein, 2018) | MIT |
| PLM/PLM-mask (Qian et al., 2021) | MIT |